Seven times some number added to 4 amounts to -20 added to the product of 3 and the number.

Answer by JP(22)

7x+4=3x-20
Subtract 3x to the left.
4x+4=-20
Subtract 4 to the right.
4x=-24
Divide -24/4 to isolate variable "x".
x=-6
Plug it back it to prove the solution.
7(-6)+4=3(-6)-20
-38=-38